OpenStudy (anonymous):

@robtobey --can you please explain this to me also 7/9 divided by 1/3 x 1/6

OpenStudy (tanya123):

\((1/3 (7/9×1/1))×1/6\)

OpenStudy (tanya123):

Answer. \(7/162\)

OpenStudy (anonymous):

tanya ur good at this stuff fan me plz?

OpenStudy (anonymous):

@tanya123 answer in book says 7/18

OpenStudy (tanya123):

Simplified answer?

OpenStudy (anonymous):

@tanya123 i did--- 7/9 x 1/2 and got 7/18 and it was right? @robtobey can you explain this better please

OpenStudy (anonymous):

Can somebody explain this,,,i thought perhaps i was understanding how to do it but i guess not, i tried the problem 5/9 divided by 1/3 x 1/6 and i got 1/3 and it was WRONG WRONG WRONG :(

OpenStudy (anonymous):

7/9 divided by 1/3 x 1/6 Refer to the Mathematica attachment.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Order_of_operations

OpenStudy (anonymous):

your great @robtobey .... Tomorrow i hope you are on... i will need some explaining tomorrow with equation similar to this \[\left( \frac{ 7 }{ 9 } \right)^2 \div \left( \frac{ 2 }{ 1 } \right) - \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

the 2/1 - 1/3 is in 1 big ( -----)

OpenStudy (anonymous):

Square the 7/9 first, then divide by 2 and finally subtract. http://en.wikipedia.org/wiki/Order_of_operations will explain the operator precedence.

OpenStudy (anonymous):

I am going to bed now,,,, I will be on tomorrow afternoon, I have 7 chapters in math left and the deadline is Aug 1st. I am NOT looking for answers just looking for expalinations that are better than the book. you've helped a bunch @robtobey

OpenStudy (anonymous):

I do have to leave now. Will check out this site tomorrow.